Lois and Clark are covering separate stories and have to travel in opposite directions. Lois leaves the Daily Planet at 8:00 AM and travels at 35 mph. Clark leaves at 8:15 and travels 40 mph. At what time will they be 140 miles apart?
Answer by checkley75(3666)

35*t+40(t-.25)=140
35t+40t-10=140
75t=140+10
75t150
t=150/75
t=2 hours after Lois leaves or @ 10:00 AM they will be 140 miles apart.
proof
35*2+40*1.75=140
70+70=140